A dairy farmer wants to mix a 75% protein supplement and a standard 25% protein ration to make 1100 pounds of a high-grade 50% protein ration. How many pounds of each should he use?
Answer by mananth(16946)

percent ---------------- quantity
Protein I 75.00% ---------------- x pounds
Protein II 25.00% ------ 1100 - x pounds
Mixture 50.00% ---------------- 1100
Total 1100 pounds
75.00% x + 25.00% ( 1100 - x ) = 50.00% * 1100
75 x + 25 ( 1100 - x ) = 55000
75 x + 27500 - 25 x = 55000
75 x - 25 x = 55000 - -27500
50 x = 27500
/ 50
x = 550 pounds 75.00% Protein I
550 pounds 25.00% Protein II

m.ananth@hotmail.ca